### v2

@@ Abstract
Free-text responses are commonly collected in psychological studies, providing rich qualitative insights that quantitative measures may not capture. Labeling curated topics of research interest in free-text data by multiple trained human coders is typically labor-intensive and time-consuming. Though large language models (LLMs) excel in language processing, LLM-assisted labeling techniques relying on closed-source LLMs cannot be directly applied to free-text data, without explicit consent for external use.
In this study, we propose a framework of assembling locally-deployable LLMs to enhance the labeling of predetermined topics in free-text data under privacy constraints. Analogous to annotation by multiple human raters, this framework leverages the heterogeneity of diverse open-source LLMs. The ensemble approach seeks a balance between the agreement and disagreement across LLMs, guided by a relevancy scoring methodology that utilizes embedding distances between topic descriptions and LLMs’ reasoning. We evaluated the ensemble approach using both publicly accessible Reddit data from eating disorder related forums, and free-text responses from eating disorder patients, both complemented by human annotations.
We found that: (1) there is heterogeneity in the performance of labeling among same-sized LLMs, with some showing low sensitivity but high precision, while others exhibit high sensitivity but low precision. (2) Compared to individual LLMs, the ensemble of LLMs achieved the highest accuracy and optimal precision-sensitivity trade-off in predicting human annotations. (3) The relevancy scores across LLMs showed greater agreement than dichotomous labels, indicating that the relevancy scoring method effectively mitigates the heterogeneity in LLMs’ labeling.
@@ 1 Introduction
In psychology, free-form text responses and qualitative data have been effectively collected since the beginning of scientific inquiry [1, 2]. To perform quantitative analyses on psychological textual data, curated topics such as symptom manifestations [3], psychosis risk factors [4], social risk factors [5], and pathological behaviors [6] must first be labeled (coded) and tailored to specific research interests. Conventionally, this process involves multiple human coders independently labeling and rating topics of interest, with reliability evaluated through the level of agreement [7]. However, this process is typically time-consuming and labor-intensive, involving both the careful training of human coders as well as their efforts in reading and labeling large volumes of text.
@@
Large language models (LLMs) are pre-trained statistical language models renowned for their ability to process and generate human-like language [8]. Research have shown that larger-sized LLMs, such as the GPTs [9, 10, 11], offer significant opportunities for facilitating data annotation [12, 13, 14, 15, 16]. However, ethical concerns regarding privacy and safety of free-text data should be central to the development of LLM-assisted methodologies [2, 17, 18]. Without explicit consent for external use, closed-source LLMs, or large-sized LLMs that cannot be deployed locally, are often unsuitable for direct application to free-text data collected in clinical or research settings [17, 18, 19].
@@
In this study, we focused on enhancing the accuracy and reliability of labeling curated topics on free-text data, using ensembles of small-sized (7-8 billion parameters), and locally-deployable LLMs, considering both security and computational constraints. Instead of inferring mental health conditions or psychological constructs, our work focused on labeling specific topics that require detailed descriptions tailored to particular research interests. The proposed ensembling framework was not designed to replace human annotation but rather to alleviate the labor burden by providing human raters with a preliminary label and relevancy score for each text.
@@
The rest of the article is organized as follows: We dedicate the remainder of the introduction (Section 1) to detailing the background of LLMs and their applications in data annotation generally, and specifically in psychology (Section 1.1). Then we discuss the rationale behind using ensembles of locally-deployable LLMs for topic labeling on free-text data (Section 1.2), followed by a related work that utilized LLM ensembles for data annotation (Section 1.3). In the methods section (Section 2), we introduce our ensemble framework (Section 2.1), and detail the evaluations using both a public and private psychological dataset (Section 2.2). Section 3 compares the ensemble with individual LLMs in terms of fidelity to human labeling. Section 4 contains a discussion of our new methodology.
1.1 LLMs for data annotation
Large language models are large-scale, pre-trained statistical language models known for their capacity and versatility in processing and generating human-like language [8]. LLMs can be categorized as open-source or closed-source based on developers’ policies [2]. Open-source models like LLaMA [20] support local deployment, enabling dataset customization and ensuring data privacy [8]. In contrast, closed-source models, such as commercial GPTs [9, 10, 11], are controlled by third parties and accessible via application programming interfaces (APIs). To adapt LLMs to specific tasks in specialized domains, two methodologies are commonly used. Fine-tuning involves further training a pre-trained model on a task-specific dataset to enhance its capabilities [21]. Prompting techniques, in contrast, adapt LLMs without additional training. The main techniques are zero-shot learning, which uses crafted prompts to direct the model’s output, and few-shot learning, which extends this by providing a few examples within the prompt, helping the model to infer the task [9].
@@ 1.1 LLMs for data annotation
Research has shown that advanced LLMs, exemplified by the GPTs, present promising opportunities for facilitating data annotation [12, 13, 14, 15, 16]. The use of larger-sized and closed-source LLMs typically outperform smaller-sized and open-source LLMs. Gilardi et al. [14] demonstrated that ChatGPT outperforms open-source LLMs in generic annotation tasks, including relevance, stance, topic, and frame detection. Alizadeh et al. [15] demonstrated that fine-tuning can enhance the performance of open-source LLMs in text classification tasks within political science, getting closer to the performance of zero-shot prompting with GPT-3.5 and GPT-4. Chew et al. [16] proposed an LLM-assisted content analysis approach using GPT-3.5, which involves iterative steps by both humans and the LLM: codebook development, labeling, and agreement evaluation, repeated until a satisfactory level of human-LLM agreement is achieved. They found that GPT-3.5 can often perform labeling at agreement levels comparable to those of human coders.
@@
In psychology, besides data augmentation [22, 23], resource enrichment [24], and developing new LLM agents for psychological applications through fine-tuning or prompting [25, 26], LLMs have been employed as classification models to label the presence or absence of mental health conditions, such as suicide and depression [27, 28], or identifying more detailed information about a mental health condition, such as severity levels and subtypes [29, 30, 2]. The usage of closed-source LLMs such as GPTs are at the forefront of these applications [2, 31]. Additionally, LLMs have been used to annotate psychological constructs. Rathje et al. [32] demonstrated that GPT-3.5-Turbo, GPT-4, and GPT-4-Turbo can accurately detect various psychological constructs (sentiment, discrete emotions, offensiveness, and moral foundations) across multiple languages using simple prompts and no additional training data, as validated by manual annotators. Peters and Matz [33] indicated that GPT-3.5 and GPT-4 can infer personality traits and psychological dispositions from social media users. Kjell et al. [19] provided evidence that, with careful validation of targeted deployment scenarios, the latest LLM technologies can transition psychological assessment from traditional rating scales to natural language communication.
@@
However, several considerations must be addressed before using LLMs to label curated psychological topics in free-text data collected from clinical or research settings. First, as a comprehensive review by Hua et al. [2] pointed out, labels of mental health conditions in existing datasets are typically not curated, with substantial variability in their definitions across studies. Curated psychological topics tailored to individual research needs—such as pathological behaviors, social or cultural factors related to a mental condition, and potential risk factors hypothesized by researchers—should be differentiated from clinical diagnoses of mental health conditions (e.g., depression or suicide) and the assessment of abstract constructs (e.g., sentiment or personality traits). Given the diversity of such curated topics and their study-specific definitions, evaluating LLMs’ performance demands a higher level of precision, complemented with human annotations tailored to each topic’s definition. In this study, we focused on curated topics with detailed definitions, ensuring that human annotations in our empirical evaluations closely aligned with the LLMs’ labeling process.
@@
Lastly, the use of a single LLM in data annotation tasks is commonplace in current applications. However, studies have shown that diverse LLMs exhibit varying capacities across tasks [35, 36], and it is reasonable to expect heterogeneity in their performance when annotating diverse topics with granular definitions. As highlighted in a survey on LLM-assisted data annotation [12], employing the Mixture of Experts concept [37] and ensembling multiple LLMs can leverage this heterogeneity to enhance overall performance and computational efficiency [38], which we elaborate on in the next section.
1.2 Ensemble of LLMs for curated topic labeling on free-text data
@@
The usage of multiple LLMs is analogous to multiple human raters who are limited to their own knowledge and biases. Studies have shown that diverse LLMs exhibit varying capacities in performing inference tasks such as coding, taxonomic labeling, and math [35, 36, satpute2024can]. We expect to observe heterogeneity in the performances of diverse open-source, same-sized LLMs when labeling various curated psychological topics. Besides differences in parameter sizes and model architectures [39], the heterogeneity in performance among LLMs also stems from the varied training datasets and preprocessing steps each LLM was exposed to, or inherited if the LLM is a variant of another [40].
@@
The main objective of ensembling diverse LLMs is to leverage the heterogeneity in LLMs and achieve reliable performance in topic labeling tasks. Ensemble learning of LLMs is typically performed either by ensembling model weights or by combining diverse outputs [41, 42]. A popular technique–LLM routing–involves training a router model using various benchmark datasets to assign LLMs to their specialized tasks and applications, aiming to balance performance and cost [43, 42, 44, 45, 46, 47]. However, this technique is not directly applicable to our task of curated topic labeling for free-text data. (1) Current router models require the inclusion of large-sized, closed-source LLMs such as GPT at the high-cost, high-performance end of the performance-cost trade-off [42, 44, 46], which is not directly applicable to free-text data. (2) The development of router models has primarily focused on assigning LLMs to their specialties in generic tasks, such as commonsense reasoning, conversation, mathematics, and code generation. Routing LLMs for the specific task of topic labeling has not been investigated, except for one study that will be discussed in detail in the related work section. The effectiveness of using this generic technique to label curated psychological topics remains unestablished. (3) Training a router model requires a pre-specified collection of LLMs and extensive training datasets, making it challenging to scale to new LLM selections, as well as diverse datasets and research needs in psychology studies.
@@
Therefore, in this study, we propose an ensembling framework that is tailored to the task of curated topic labeling for free-text data. This framework is scalable in LLM selections and can be readily adapted to individual research needs without requiring the training of a new model. The framework leverages the heterogeneity of diverse, open-source, and locally-deployable LLMs to enhance the accuracy and reliability of determining the presence or absence of a topic in a given text. The ensembling seeks a balance between the agreement and disagreement across multiple LLMs, guided by a novel relevancy scoring methodology. We evaluated the ensembling framework using both a public dataset related to a mental health condition and a protected clinical dataset, both complemented by human annotations. For the public dataset, we also employed the large-sized GPT-4o as a judge [48]. This framework was not designed to replace human evaluation but rather to alleviate the labor burden by providing human raters with a preliminary label and relevancy score for each text.
1.3 Related work
@@
A very recent work utilized the ensemble of LLMs for data annotation tasks [49]. Farr et al. [49] proposed an ensembling methodology that aligns multiple LLMs in a chain, routing subsets of data to subsequent models based on the classification uncertainty measured for the current model. This ensemble is an extension of the LLM routing technique but does not require to train a router model. It was shown to outperform individual models in the chain and reduce the use of high-cost LLMs. However, our ensembling framework differs significantly from theirs. First, as an extension of the LLM routing technique, their designs incorporated both open-source smaller-sized LLMs and a large-sized, closed-source GPT-4o. The extent to which performance improvements attributed to the use of GPT-4o remains unclear. Given the constraints of free-text data, our method mainly focuses on enhancing the performance of open-source deployable LLMs through ensembling. Second, their method suggests that more robust models be placed later in the chain, with the robustness indicated by a progression from smaller to larger model sizes. In our study, we aim to evaluate the advancements of ensembling compared to individual LLMs, controlling for the same model sizes. In addition to the F1-score assessed in their experiment, we evaluated the performance through multiple lenses, including precision, sensitivity and inter-rater reliability. Lastly, the labeling tasks evaluated in their empirical experiments—stance detection, misinformation detection, and ideology detection—are abstract components of text, that do not directly align with the needs for precise labeling of curated topics in psychological studies.

### v4

@@ Abstract
Free-text responses are commonly collected in psychological studies, providing rich qualitative insights that quantitative measures may not capture. Labeling curated topics of research interest in free-text data by multiple trained human raters is typically labor-intensive and time-consuming. Though large language models (LLMs) excel in language-based tasks, LLM-assisted labeling techniques relying on closed-source LLMs cannot be directly applied to clinical free-text data, without explicit consent for external use.
Analogous to labeling by multiple human raters, we propose a framework of assembling locally-deployable LLMs to improve the process of the labeling of predetermined topics in free-text data under privacy constraints. This framework leverages the heterogeneity of diverse open-source LLMs, and the ensemble approach seeks a balance between the agreement and disagreement across LLMs, guided by a relevancy scoring methodology that utilizes embedding distances between topic descriptions and LLMs’ reasoning. We evaluated the ensemble approach using both publicly accessible Reddit data from eating disorder related forums, and free-text responses from eating disorder patients, both complemented by human annotations.
We found that: (1) there is heterogeneity in the performance of labeling among same-sized LLMs, with some showing low sensitivity but high precision, while others exhibit high sensitivity but low precision. (2) Compared to individual LLMs, the ensemble of LLMs achieved the highest accuracy and optimal precision-sensitivity trade-off, as evaluated by human annotations. (3) The relevancy scores across LLMs showed greater agreement than dichotomous labels, indicating that the relevancy scoring method effectively mitigates the heterogeneity in LLMs’ labeling.
@@ 1 Introduction
In psychology, free-form text responses and qualitative data have been effectively collected since the beginning of scientific inquiry [1, 2]. To perform quantitative analyses on psychological textual data, curated topics such as symptom manifestations [3], psychosis risk factors [4], social risk factors [5], and pathological behaviors [6] must first be labeled (coded) and tailored to specific research interests. Conventionally, this process involves multiple human coders independently labeling topics of interest, with reliability evaluated through the level of agreement [7]. However, this process is typically time-consuming and labor-intensive, involving both the careful training of human coders as well as their efforts in reading and labeling large volumes of text.
@@
Pre-trained large language models (LLMs) are statistical models that can be adapted for a variety of downstream tasks. Research have shown that larger-sized LLMs, such as the GPTs [8, 9, 10], offer significant opportunities in data annotation tasks [11, 12, 13, 14, 15]. However, ethical concerns regarding privacy and safety of free-text data should be central to the development of LLM-assisted methodologies [2, 16, 17]. Without explicit consent for external use, closed-source LLMs, or large-sized LLMs that cannot be deployed locally, are often unsuitable for direct application to free-text data collected in clinical or psychological research settings [16, 17, 18].
@@
In this study, we focus on enhancing the accuracy and reliability of labeling curated topics on free-text data, while accounting for both security and computational constraints in clinical and psychological research. Instead of inferring psychological conditions or constructs, our work focus on labeling specific topics that require detailed descriptions tailored to individual research interests. Analogous to the labeling process by multiple human raters, we propose a framework that assembles multiple small-sized (7–8 billion parameter), locally deployable language models to label and score the presence of target topics in free-text data.
@@
The rest of the article is organized as follows: We dedicate the remainder of the introduction to detailing the background of LLMs (1.1) and their applications in data annotation generally, and specifically in psychology (Section 1.2). We then discuss the common ensemble technique and related work to motivate the use of locally deployable LLM ensembles for labeling target topics in free-text data (Section 1.3). In the methods section (Section 2), we introduce our ensemble framework (Section 2.1), and detail the evaluations using both a public and private psychological dataset (Section 2.2). Section 3 compares the ensemble with individual LLMs in terms of fidelity to human labeling. Section 4 contains a discussion of our proposed framework.
1.1 Background of Large Language Models
Large language models are large-scale, pre-trained neural networks known for their capacity and versatility in processing and generating human-like language [19]. Popular LLMs typically rely on the Transformer architecture and its attention mechanism [20], which helps to encode not only the semantic meaning of words but also their relationships and surrounding context. Natural language inputs are first transformed into continuous vectors that encode semantic and positional information. These embeddings are then typically processed by an encoder–decoder (sequence‑to‑sequence) [21, 20, 22] or a decoder‑only (autoregressive) architecture [23, 24, 25] to generate the most probable next token. LLMs can be adapted for a variety of downstream tasks including text classification [26, 27, 28, 29], topic modeling [30, 31], and question answering [32, 33, 34, 35]. For language-based usages, LLMs can be broadly categorized into models designed for representation learning and models designed for generation. Representation-focused models, such as encoder-only or encoder–decoder architectures (e.g., BERT [21], RoBERTa [36]), are primarily trained to produce contextual embeddings for downstream tasks like classification or retrieval. In contrast, generative models, typically decoder-only architectures (e.g., GPTs [8, 9, 10]), are trained to autoregressively generate coherent and human-like language responses.
Importantly, with respect to accessibility and security, LLMs can be categorized as open-source or closed-source depending on developers’ policies [2]. Open-source models, such as LLaMA [37], support local deployment, enabling dataset customization and ensuring data privacy [19]. In contrast, closed-source models, such as commercial GPTs [8, 9, 10], are controlled by third parties and accessed through application programming interfaces (APIs).
1.2 Language Models for Data Annotation
@@
Research has shown that advanced LLMs present promising opportunities for facilitating data annotation [11, 12, 13, 14, 15]. The use of larger-sized and closed-source LLMs typically outperform smaller-sized and open-source LLMs. Gilardi et al. [13] demonstrated that ChatGPT outperforms open-source LLMs in generic annotation tasks, including relevance, stance, topic, and frame detection. Alizadeh et al. [14] demonstrated that fine-tuning can enhance the performance of open-source LLMs in text classification tasks within political science, getting closer to the performance of zero-shot prompting with GPT-3.5 and GPT-4. Chew et al. [15] proposed an LLM-assisted content analysis approach using GPT-3.5, which involves iterative steps by both humans and the LLM: codebook development, labeling, and agreement evaluation, repeated until a satisfactory level of human-LLM agreement is achieved. They found that GPT-3.5 can often perform labeling at agreement levels comparable to those of human coders.
@@
In psychology, LLMs have also been employed to annotate psychological content, along with tasks such as data augmentation [38, 39], resource enrichment [40], and the development of new LLM agents [41, 42]. Specifically, LLMs have been used as classification models to label the presence or absence of general mental health conditions (e.g., suicide and depression [43, 44]), as well as to provide more detailed information about these conditions (e.g., severity levels and subtypes [45, 46, 2]). Notably, closed-source LLMs such as GPTs are at the forefront of these applications [2, 47]. Additionally, closed-source LLMs have been employed to annotate psychological constructs. Rathje et al. [48] demonstrated that GPT-3.5 and GPT-4 can accurately identify constructs such as sentiment, discrete emotions, and moral foundations across multiple languages using simple prompts without additional training, with performance validated by human annotators. Similarly, Peters and Matz [49] showed that GPT-3.5 and GPT-4 can infer personality traits and psychological dispositions from social media data. Extending further, Kjell et al. [18] provided evidence that, with careful validation of deployment scenarios, the latest LLM technologies can shift psychological assessment from traditional rating scales toward natural language communication.
However, several considerations must be addressed before using LLMs to label curated psychological topics in free-text data collected from clinical or research settings.
@@
First, as a comprehensive review by Hua et al. [2] pointed out, labels of mental health conditions in existing datasets are typically not curated, with substantial variability in their definitions across studies. Curated psychological topics tailored to individual research needs—such as pathological behaviors, social or cultural factors related to a mental condition in a certain population, and potential risk factors hypothesized by researchers—should be differentiated from clinical diagnoses of mental health conditions (e.g., depression or suicide) and the assessment of abstract constructs (e.g., sentiment or personality traits). Given the diversity of such curated topics and their study-specific definitions, evaluating LLMs’ performance demands a higher level of precision, complemented with human annotations tailored to each topic’s definition. In this study, we focused on curated topics with detailed definitions, ensuring that human annotations in our empirical evaluations closely aligned with the LLMs’ labeling process.
@@
Lastly, the use of a single LLM in data annotation tasks is commonplace in current applications. However, studies have shown that diverse LLMs exhibit varying capacities across tasks [51, 52], and we should expect heterogeneity in their performance when annotating diverse topics with granular definitions. As highlighted in a survey on LLM-assisted data annotation [11], employing the Mixture of Experts concept [53] and ensembling multiple LLMs can leverage this heterogeneity to enhance overall performance and computational efficiency [54], which we elaborate on in the next section.
1.3 LLM Ensembles for Labeling Curated Topics in Free-text Data
@@
Studies have shown that diverse LLMs exhibit varying capacities in performing inference tasks such as coding, taxonomic labeling, and math [51, 52, 55]. We expect to observe heterogeneity in the performances of diverse open-source, same-sized LLMs when labeling various curated psychological topics. Besides differences in parameter sizes and model architectures [56], the heterogeneity in performance among LLMs also stems from the varied training datasets and preprocessing steps each LLM was exposed to, or inherited if the LLM is a variant of another [57].
@@
The main objective of ensembling diverse LLMs is to leverage the heterogeneity in LLMs and achieve reliable performance in topic labeling tasks. Ensemble learning of LLMs is typically performed either by ensembling model weights or by combining diverse outputs [58, 59]. A common technique–LLM routing–involves training a router model using various benchmark datasets to assign LLMs to their specialized tasks and applications, aiming to balance performance and cost [60, 59, 61, 62, 63, 64]. However, this technique is not directly applicable to our task of curated topic labeling for free-text data. (1) Current router models require the inclusion of large-sized, closed-source LLMs such as GPT at the high-cost, high-performance end of the performance-cost trade-off [59, 61, 63], which is not directly applicable to free-text data. (2) The development of router models has primarily focused on assigning LLMs to their specialties in generic tasks, such as commonsense reasoning, conversation, mathematics, and code generation. Routing LLMs for the specific task of topic labeling has not been investigated, except for one study that will be discussed in detail in the related work section. The effectiveness of using this generic technique to label curated psychological topics remains unestablished. (3) Training a router model requires a pre-specified collection of LLMs and extensive training datasets, making it challenging to scale to new LLM selections, as well as diverse datasets and research needs in psychology studies.
@@
Therefore, in this study, we propose an ensembling framework tailored to the task of curated topic labeling in free-text data. Our use of multiple LLMs is analogous to employing multiple human raters, who are trained to label curated topics for specific research questions but are nevertheless constrained by their individual knowledge and biases. Importantly, we focus on the task of identifying predetermined topics of interest (a top-down approach), which is common in human annotation, rather than a data-driven exploratory topic analysis task using methods such as Latent Dirichlet Allocation (a bottom-up approach). Our framework is scalable in LLM selections and can be readily adapted to individual research needs without requiring the training of new models. The framework leverages the heterogeneity of diverse, open-source, and locally-deployable LLMs to enhance the accuracy and reliability of determining the presence or absence of a topic in a given text. The ensembling seeks a balance between the agreement and disagreement across multiple LLMs, guided by a relevancy scoring methodology. We evaluated the ensembling framework using both a public dataset related to a mental health condition and a protected clinical dataset, both complemented by human annotations. For the public dataset, we also employed the large-sized GPT-4o as a judge [65]. This framework was not designed to replace human evaluation but rather to alleviate the labor burden by providing human raters with a preliminary label and relevancy score for each text. Finally, this framework is uniquely accessible to researchers with fewer computational resources, as the LLMs being assembled can be run on consumer grade computers.
@@
Another recent work utilized the ensemble of LLMs for data annotation tasks without training models, but our ensembling framework differs significantly from theirs. Farr et al. [66] proposed an ensembling methodology that aligns multiple LLMs in a chain, routing subsets of data to subsequent models based on the classification uncertainty measured for the current model. This ensemble is an extension of the LLM routing technique but does not require to train a router model. It was shown to outperform individual models in the chain and reduce the use of high-cost LLMs. Our ensembling framework differs from theirs in the following ways. First, as an extension of the LLM routing technique, their designs incorporated both open-source smaller-sized LLMs and a large-sized, closed-source GPT-4o. The extent to which performance improvements attributed to the use of GPT-4o remains unclear. Given the constraints of free-text data, our method mainly focuses on enhancing the performance of open-source deployable LLMs through ensembling. Second, their method suggests that more robust models be placed later in the chain, with the robustness indicated by a progression from smaller to larger model sizes. In our study, we aim to evaluate the advancements of ensembling compared to individual LLMs, controlling for the same model sizes. In addition to the F1-score assessed in their experiment, we evaluated the performance through multiple lenses, including precision, sensitivity and inter-rater reliability. Lastly, the labeling tasks evaluated in their empirical experiments—stance detection, misinformation detection, and ideology detection—are abstract components of text, that do not directly align with the needs for precise labeling of curated topics in psychological studies.
Figure 1: The ensemble framework of multiple LLMs (right) is analogous to the labeling process of multiple human raters (left) in the task of rating curated topics in free-text data. It leverages the heterogeneous performance of diverse open-source, deployable language models. (1) Labeling and scoring: Each model produces a label and a score for a given topic in an input text. Specifically, each model returns a dichotomous label indicating the presence of the topic, along with text phrases related to the topic that reflect its reasoning. A relevance score is then computed to measure the semantic similarity between the identified phrases and the predefined topic description. This is analogous to a human rater deciding whether a topic is present in a text (e.g., a patient narrative) and providing the phrases they identified as indicators. (2) Ensemble across LLMs: For a given topic, an ensemble score is calculated by averaging model scores (e.g., via PCA). The ensemble label is then determined by removing false positives from the union of positive cases across LLMs (excluding outliers), guided by the ensemble score. This mirrors majority voting combined with group reasoning among human raters to reach a final consensus.